Solve by elimination method
2x+3y=-5
5x-3y=-23

Answer by mananth(16949)

2x+3y=-5
5x-3y=-23
..
Add the 2 equations to eliminate y
2x+3y+5x-3y=-5-23
7x=-28
divide by 7
7x/7 = -28/7
x=-4
..
plug the value of x in any of the equations
2x+3y=-5
2*-4+3y=-5
-8+3y=-5
Add 8 to both sides
-8+8+3y=-5+8
3y=3
divide by 3
3y/3 = 3/3
y=1